A random sample of 376 married couples found that 298 had two or more personality preferences in common. In another random sample of 562 married couples, it was found that only 26 had no preferences in common. Let p1 be the population proportion of all married couples who have two or more personality preferences in common. Let p2 be the population proportion of all married couples who have no personality preferences in common.

(a) Find a 99% confidence interval for p1p2. (Use 3 decimal places.)

lower limit    
upper limit    
Answer #1

p1cap = 298/376 = 0.7926
p2cap = 26/562 = 0.0463

pcap = (298 + 26)/(376 + 562) = 0.3454

SE = sqrt(0.3454*(1-0.3454)*(1/376 + 1/562))
SE = 0.0317

lower limit = (0.7926 - 0.0463 - 2.5758*0.0317) = 0.665

upper limit = (0.7926 - 0.0463 + 2.5758*0.0317) = 0.8280